Solve the system of equation
5x+y=12
2x=3y-19

Answer by checkley75(3666)

5X+Y=12 OR Y=-5X+12 NOW REPLACE Y IN THE OTHER EQUATION WITH (-5X+12)
2X=3(-5X+12)-19
2X=-15X+36-19
2X+15X=17
17X=17
X=1 ANSWER.
5*1+Y=12
5+Y=12
Y=12-5
Y=7 ANSWER.
PROOF
2*1=3*7-19
2=21-19
2=2